A cart with a mass of 5.0 kg accelerates from rest due to a force of 10.0 N being applied to it at an angle of 30 degrees above the horizontal. Assuming no friction, what is the velocity of the cart after 5 seconds?




1
Expert's answer
2021-02-22T10:24:51-0500
"v=at\\v=\frac{F\cos{30}}{m}t\\v=\frac{10\cos{30}}{5}5=8.66\frac{m}{s}"